Solve the system using the substitution method. (If there is no solution, enter NO SOLUTION.)
{3x + 2y = −1
{x= y + 13

Answer by Alan3354(69443)

3x + 2y = -1
x = y + 13
---------
Sub for x in the 1st equation
3(y+13) + 2y = -1
Solve for y.
Then solve for x.